OpenStudy (anonymous):

Solve for x: Two thirds (x − 4) = 2x

OpenStudy (anonymous):

(2/3)(x-4)=2x ?

OpenStudy (anonymous):

(2/3)(x-4)=2x Divide 2/3 on both sides x-4=3x Move x on one side x-4-x=3x-x -4=2x Divide both side by 2 -4/2=2x/2 x=-2